Two moles of helium, four moles of hydrogen and one mole of water vapour form an ideal gas mixture. The molar specific heat at constant pressure of the mixture is(neglecting vibrational degrees of freedom):

a> (7/3)R

B> (16/7)R

C> (10/3)R

D> (23/7)R


.......................................

THE CORRECT ANSWER IS (23/7) R

PLEASE GIVE ME A DETAILED SOLUTION


Grade:9

2 Answers

SURA CHAITHANYA
6 years ago
Cv = 4*5R/2 + 2*3R/2 +1*6R/2 /7    =16R/7
Cp = Cv + R = 16R/7 + R  =23R/7
Cv = n1Cv1 + n2Cv2 + n3Cv3  /n1 + n2 + n3
ankit singh
3 years ago

CV for hydrogen =  25R
Cv for helium =  23R
VV for water vapour =  26R=3R
(CV)mix =  4+2+14× 25R+2× 23R+1×3R =  716R
CP = CV+R
CP =  716R+R
Or, CP =  723R